How large a mass will 32 N accelerate at the same rate achieved by 13 N acting one 8.3 kg?
1
Expert's answer
2019-02-06T12:52:03-0500

Let's first find the acceleration produced by

"13 N"

force acting on

"8.3 kg"

mass:

"a = \frac{F}{m} = \frac{13 N}{8.3 kg} = 1.56 \frac{m}{s^2}."

Finally, we can find the mass that will accelerate by

"32 N"

force acting at the same rate achieved by 13 N acting one 8.3 kg:

"m = \frac{F}{a} = \frac{32 N}{1.56 \frac{m}{s^2}} = 20.5 kg."

Answer:

"m = 20.5 kg"

.